# Era Splitting - Invariant Learning for Decision Trees

### v4

@@ 1 Introduction
Since the base learner $f$ is a decision tree, optimizing the decision tree function is equivalent to finding the optimal split point of the input data $x_{i}$ at each step. Each split point is defined by a single feature and a single value of that feature’s data. Starting at the the root node, which contains all the training data, the optimal split point is chosen. All the data with that feature greater than or equal to the split value gets partitioned into the left child node, and the data points with the feature less than the split value go to the right child node. This is a decision tree of depth 1. In successive steps the child nodes become parent nodes, and the data in those nodes is split in the same manner as the root node. The function that determines the optimal split point at each node is called the splitting criterion, which measures the reduction in impurity (also called information gain) induced by the split. The splitting criterion assigns a score to each possible split of the data at each node. The split is chosen which has the highest score, as this is the optimal split, resulting in the purest child nodes. Impurity is measured in the targets, or labels. Good splits create child nodes with the low impurity. For an introduction to decision trees and a review of the popular splitting criteria and impurity measures, refer to the text ([39]). In GBDTs for regression, the popular spitting criterion implements a custom measure of impurity described in ([26]), $\mathcal{L}_{\text{imp}}$, in equation 3.
$\mathcal{L}_{\text{imp}}=-\frac{1}{2}\sum_{j=1}^{T}\frac{\left(\sum_{i\in I_{j}}g_{i}\right)^{2}}{\sum_{i\in I_{j}}h_{i}+\lambda}-\gamma T$ (3)
$\mathcal{L}_{\text{imp}}$ computes the quality of an entire tree structure, where $g_{i}$ and $h_{i}$ are the gradients and hessians in leaf node $j\in\{1,...,T\}$ of the convex differentiable loss function. It is analogous to computing the variance in the node, but this function attains a maximum for minimum variance. In practice it isn’t feasible to evaluate every possible tree structure, but instead the iterative process grows a tree from the root node by selecting the best split based on the decrease in impurity from the parent to the child nodes.
$\mathcal{L}_{\text{split}}=\frac{1}{2}\left[\frac{\left(\sum_{i\in I_{L}}g_{i}\right)^{2}}{\sum_{i\in I_{L}}h_{i}+\lambda}+\frac{\left(\sum_{i\in I_{R}}g_{i}\right)^{2}}{\sum_{i\in I_{R}}h_{i}+\lambda}-\frac{\left(\sum_{i\in I}g_{i}\right)^{2}}{\sum_{i\in I}h_{i}+\lambda}\right]-\gamma$ (4)
Equation 4 is the original splitting criterion. $I$, $I_{L}$, and $I_{R}$ are data set identifiers corresponding to the data of the parent node, the left child node, and the right child node respectively. While equation 3 sums over all $T$ leaf nodes, the splitting criterion only evaluates the impurity of the current child nodes compared to the parent node. For GBDTs, implementing regression with the mean-squared error loss, a common simplification is that the Hessians are constant and equal to 1. Highest scores imply the greatest reduction in variance in the child nodes.
@@
In the basic setting, GBDTs pool all the training data together when computing the splitting criterion. If the training data inherently comes from separate environments (eras), as defined in the OOD literature, what is lost by pooling all the training data together to find the best split? Would it be better to employ a splitting criterion that treats data from each environment (era) separately? This research endeavors to show that the answer to this question is, ”yes”.
@@
A new splitting criterion called era splitting is proposed in section 2. It combines the impurity reduction calculation from each era separately into one final new criterion, which is a smooth average of the impurity reduction of the era-wise splits. It captures the essence of invariance from the IRM paradigm applied to tree splitting criteria. An additional era-wise splitting criterion is developed in section 3, which is designed after the AND-mask, and looks for splits which maximize the agreement in the direction of the predictions implied by the split for each era (environment). In section 4, the theoretical interpretation is expounded briefly, giving three new propositions. Section 5 then describes the experimental design used to validate our new splitting criteria. The experiments include a one-dimesnional toy model aimed at verifying the objectives of our new splitting, a high-dimensional synthetic dataset proposed for the same purpose in ([37]), and two real-world empirical studies. The first is the Camelyon17 dataset ([23]) breast cancer detection domain generalization problem, the second is financial stock market cross-sectional return problem provided by the Numerai data science tournament. In section 6, the results from the experiments are presented. Finally section 7 provides interpretation of the results and discussion about future work.
@@ 2 Era Splitting
The term era is used in the same sense as the term environment is used in the references ([21, 37]) and sometimes the term domain as in ([33]). An era of data is a disjoint set by which our training and test data sets are comprised of many eras of data. See definition 4.1 below. The general understanding is that each era contains data that occurs at certain distinct times and/or physical locations. In the context of the splitting criterion (4), the data set definitions $I$, $I_{L}$, and $I_{R}$ need to be refined when dealing with environments. With era splitting, each training data point comes from one of $M$ distinct eras. Thus there exists data sets $I^{j}$, $I_{L}^{j}$, and $I_{R}^{j}$, referring to data coming from the $j^{th}$ era, such that $I^{j}\in I$, $I_{L}^{j}\in I_{L}$, and $I_{R}^{j}\in I_{R}$ for $j\in\{1,2,...,M\}$. Using this convention, a new era-wise splitting criterion, which computes the information gain of a split for the $j^{th}$ era of data is given.
$\mathcal{L}_{\text{split}}^{j}=\frac{1}{2}\left[\frac{\left(\sum_{i\in I_{L}^{j}}g_{i}\right)^{2}}{\sum_{i\in I_{L}^{j}}h_{i}+\lambda}+\frac{\left(\sum_{i\in I_{R}^{j}}g_{i}\right)^{2}}{\sum_{i\in I_{R}^{j}}h_{i}+\lambda}-\frac{\left(\sum_{i\in I^{j}}g_{i}\right)^{2}}{\sum_{i\in I^{j}}h_{i}+\lambda}\right]-\gamma$ (5)
The era splitting criterion in its basic form is written as a mean average of the era-wise splitting criterion for each era of data.
@@
The era splitting criterion will have the highest score for the split which provides the greatest average decrease in impurity over all of the eras of data. Let us notice that this new criterion recovers equation 4 in the case where there is only one era of data. Therefore the era splitting criterion can be viewed as a generalization of the original splitting criterion. This criterion will be less-likely to favor splits which work really well in some eras but not in others. Splits are not considered which are undefined in any era. Potential split points must have data to the left and right of them to be considered as optimal. The reason for this choice is illustrated in figure 3. The plot on the left shows a data configuration where the original splitting criterion will choose a split which doesn’t induce any impurity decrease in any era of data. On the other era hand, era splitting chooses a split reducing impurity in both eras simultaneously.
A further generalization is made by way of a smooth max function, called the Boltzmann operator ([22]). The Boltzmann operator operates on an array of real numbers, and has one parameter, called $\alpha$. When $\alpha=0$, the Boltzmann operator recovers the arithmetic mean, when $\alpha=-\infty$ it recovers the $\min$, and when $\alpha=\infty$ it recovers the $\max$. To define the Boltzmann operator ($\mathcal{B}_{\alpha}$), consider $n$ real numbers, $x_{i}$ for $i\in\{1,...,n\}$.
$\mathcal{B}_{\alpha}(x_{1},...,x_{n})=\frac{\sum_{i=1}^{n}x_{i}\exp{(\alpha x_{i})}}{\sum_{i=1}^{n}\exp{(\alpha x_{i})}}$ (7)
The final version of the era splitting criterion is defined in equation 8, which is understood as the smooth maximum (minimum) of the era-wise impurity reduction over all of the eras, and is a generalization of the original splitting criterion in Eq 4.
$\mathcal{L}_{\text{era split}}^{\alpha}=\mathcal{B}_{\alpha}\left(\mathcal{L}_{\text{split}}^{1},...,\mathcal{L}_{\text{split}}^{M}\right)$ (8)
@@ 3 Directional Era Splitting
There is a subtle ambiguity that arises with equation 8, which requires the introduction of the concept of the direction implied by a split. A split implies a direction via the values of the child nodes. See figure 2. If the value of the left child node is larger than the value of the right child node, then the direction of this split is too the left. If the value of the right child node is larger, then the direction is to the right. Let the value of the left child node be $v_{l}$, and the value of the right child node be $v_{r}$, then without loss of generality, the direction of the split, $d_{\text{split}}$, is defined as the sign of the difference between the left child node value and the right child node value.
$d_{\text{split}}=\text{sign}(v_{l}-v_{r})$ (9)
In the traditional setting where all the training data is pooled together, there can be only one direction implied over the entire data set. The traditional splitting criterion (eq. 4) does not consider the direction of the split, since it is superfluous. It doesn’t matter whether the direction implied by the split is to the left or to the right, since there is only one era, just as long as the impurity is reduced. In era splitting, the split criterion is computed for each era separately, and then summarized with (the Boltzmann operator). A problem that arises here is that there could be splits that imply conflicting directions from one era to the next. We can describe the direction of a split in era $j$ by adapting equation 9 with era-wise descriptors, as follows.
@@
The directions implied by the split for each era of data results in an array of length $M$, $\{d_{\text{era split}}^{0},d_{\text{era split}}^{1},...,d_{\text{era split}}^{M}\}$. Directional era splitting, $\mathcal{D}_{\text{era split}}$, prefers splits which have the most agreement in the direction across all eras.
$\mathcal{D}_{\text{era split}}=\frac{1}{M}\left|\sum_{j=0}^{M}d_{\text{era split}}^{j}\right|$ (11)
Each $d_{\text{era split}}^{j}$ will take a value of either $1$ or $-1$, depending on the direction of the split for that era. The sum is bounded on the interval $[-M,M]$. By taking the absolute value and dividing my $M$, $\mathcal{D}_{\text{era split}}$ is then bounded on the interval $[0,1]$. At the low end, with a value of 00, half of the directions go in one way while half go in the other way, the highest level of disagreement. At the high end, with a value of $1$, all the directions go the same way, the most agreement. When choosing the best split during tree growth, higher values of $\mathcal{D}_{\text{era split}}$ are preferred.
@@ 4 Theoretical Breakdown
This section expounds three propositions which are illustrative of the motivations behind era splitting. Here the focus is on era splitting (eq. 8), and not directional era splitting (eq. 11). The former being more anchored in traditional theory, and connected with the original criterion. The later is new and experimental, with impurity tied to the direction of the predictions in each era. Nevertheless, the following proposition 4.3 applies to both.
The problem with traditional splitting methods in the OOD setting is that the ”optimal” split can be degenerate in any or all of the eras (environments). The term, degenerate, indicates that the split doesn’t result in improved impurity in the child nodes compared to the parent for any particular era of the data set. This is a split that never would have been if the tree was grown on that era of data independently. Eras ( or environments ) usually occur independently in nature, such as disjoint time frames or locations. The left side of figure 3 shows a simple example of a degenerate split with a toy data set of 4 data points. The original split criterion favors the split (red dashed line) which effectively separates era 1 data from era 2 data. In each era, this split does not induce any reduction in impurity. The era splitting criterion, on the other hand, favors a split (green dotted line) which reduces impurity in both eras simultaneously.
Definition 4.1 (Era (Environment)).
Given the input-output training data $(X,Y)$ consisting of $N$ data points, an era (environment) of the data is a subset of the entire training set, $(X,Y)_{j}$ for $j\in\{1,..,M\}$, for which there are $M$ eras such that $\bigcup_{j}^{M}(X,Y)_{j}=(X,Y)$, and the eras do not intersect.
Definition 4.2 (Degenerate Splits).
A degenerate split is one which does not reduce impurity in any era of a training data.
Proposition 4.1.
@@
Table 1: Data for proposition 4.1.
@@
This is one of the main conceptual motivations for the current research: it is possible that the original method misses, ignores, or misinterprets key environmental information. Era splitting addresses this problem in a very direct way: evaluate each split based on each environment separately. There is one configuration of era splitting which ensures the split points will always reduce impurity in every environment (era) of data simultaneously, never producing degenerate splits. This result is presented in proposition 4.2.
Proposition 4.2.
The era splitting criterion as defined in equation 8 with Boltzmann alpha parameter set to $\infty$- ∞ will never induce a degenerate split.

### v5

@@ 1.3 Splitting
Optimizing a regression tree $f$ is equivalent to finding the optimal split points of the input data $x_{i}$ at each node until a terminal node is reached. Each split point is defined by a single feature and a single value of that feature’s data. All the data with that feature less than or equal to the split value gets partitioned into the left child node, and the data points with the feature greater than the split value go to the right child node. In successive steps the child nodes become parent nodes, and the data in those nodes is split in the same manner as the root node. Nodes are split like this until terminal nodes are reached. The function that determines the optimal split point at each node is called the splitting criterion, which measures the reduction in impurity. The split criterion was introduced in equation (7). The minimum value of $\Delta R(s,t)$ over all splits $s$ in the set of all splits $\mathcal{S}$ is the split which most reduces most the average error of the tree prediction, $R(T)$, as per definition 2 above. Minimizing $\Delta R(s,t)$ is equivalent to maximizing the original split criterion, given in the next definition.
Definition 3 (Original Split Criterion, Chen and Guestrin [2016b]).
Define the original split criterion as
$\mathcal{L}_{\text{split}}=\frac{1}{2}\left[\frac{\left(\sum_{i\in I_{L}}g_{i}\right)^{2}}{\sum_{i\in I_{L}}h_{i}+\lambda}+\frac{\left(\sum_{i\in I_{R}}g_{i}\right)^{2}}{\sum_{i\in I_{R}}h_{i}+\lambda}-\frac{\left(\sum_{i\in I}g_{i}\right)^{2}}{\sum_{i\in I}h_{i}+\lambda}\right],$ (26)
where $I$, $I_{L}$, and $I_{R}$ are data set identifiers corresponding to the data of the parent node, the left child node, and the right child node respectively. The $g_{i}$ are the gradients as defined in algorithm 2, above, $h_{i}$ are the hessians (which are constant for regression), and $\lambda$ is the L2-regularization term.
With the original split criterion defined, we have outlined the state-of-the-art in GBDTs, which we will take as baseline. The original contributions of this paper will be 2 new alternative splitting criteria inspired by recent trends in the OOD generalization research. For this reason we need to first define eras more concisely.
1.4 Eras (Environments)
This research employs existing conventions to define what are called eras, also known as environments Peters et al. [2015], Arjovsky et al. [2017] and sometimes domains as in Koh et al. [2021]. This paragraph paraphrases from Peters et al. [2015], which is the standard setup in related works Arjovsky et al. [2020] and Parascandolo et al. [2020]. Firstly, assume there are different experimental conditions belonging to a the set of all experimental conditions, $e\in\mathcal{E}$, and we take an i.i.d. sample $(\mathrm{X}^{e},\mathrm{Y}^{e})\in(\mathbb{R}^{d},\mathbb{R})$ from each environment, where $\mathrm{X}^{e}$ and $\mathrm{Y}^{e}$ are the independent and dependent variables respectively. In particular, the different distributions of $X^{e}$ in the environments are unknown and not precisely controlled.
Definition 4 (Era (Environment) Peters et al. [2015]).
An era (environment) refers to a particular experimental conditions $e$ from the set of all possible conditions $\mathcal{E}$.
For this research, it is assumed that every data point originates from some environment, and our entire data set is the union of the data from all the environments. Practically speaking, this comes down to assigning an environmental identifier $j$ to each data point. The identifier is an integer referring to the index of the environment the data comes from. In the following text it is usually assumed that we have data from $M$ eras, and so the training data $(X,Y)$ is the union of data from all $M$ environments
$(X,Y)=\bigcup_{j=1}^{M}(X^{j},Y^{j}).$ (27)
@@ 1.4 Eras (Environments)
In the original setting, GBDTs pool all the training data together when computing the splitting criterion. If the training data inherently comes from separate environments (eras), as defined in the OOD literature Peters et al. [2015], Arjovsky et al. [2020], what is lost by pooling all the training data together to find the best split? Would it be better to employ a splitting criterion that treats data from each environment (era) separately? This research endeavors to show that the answer to this question is, ”yes”.
1.5 New Contributions and Outline
@@
A new splitting criterion called era splitting is proposed in section 2. It combines the impurity reduction calculation from each era separately into one final new criterion, which is a smooth average of the impurity reduction of the era-wise splits. It captures the essence of invariance from the IRM paradigm Arjovsky et al. [2020] applied to the original split criterion. Another new splitting criterion, called directional era splitting, is presented in section 3, which is designed after the AND-mask Parascandolo et al. [2020], and looks for splits which maximize the agreement in the direction of the predictions implied by the split for each era (environment). In section 4, some theoretical concepts are discussed, highlighting the motivation for the new splitting criteria. Section 5 then describes the experimental design used to validate our new splitting criteria. The experiments include a one-dimensional toy model aimed at verifying the objectives in a simple setting, a high-dimensional synthetic dataset proposed for the same purpose in Parascandolo et al. [2020], and two real-world empirical studies. The first is the Camelyon17 dataset Bandi et al. [2018] breast cancer detection domain generalization problem, the second is the financial stock market cross-sectional return problem provided by the Numerai data science tournament. In section 6, the results from the experiments are presented. Finally section 7 provides interpretation of the results and discussion about future work.
@@ 2 Era Splitting
In the context of the original splitting criterion in equation (26), the data set definitions $I$, $I_{L}$, and $I_{R}$ need to be refined when dealing with environments. With era splitting, each training data point comes from one of $M$ distinct eras. Thus there exists data sets $I^{j}$, $I_{L}^{j}$, and $I_{R}^{j}$, referring to data coming from the $j^{th}$ era, such that $I^{j}\in I$, $I_{L}^{j}\in I_{L}$, and $I_{R}^{j}\in I_{R}$ for $j\in\{1,2,...,M\}$. Using this convention, a new era-wise splitting criterion, which computes the information gain of a split for the $j^{th}$ era of data is given by
$\mathcal{L}_{\text{split}}^{j}=\frac{1}{2}\left[\frac{\left(\sum_{i\in I_{L}^{j}}g_{i}\right)^{2}}{\sum_{i\in I_{L}^{j}}h_{i}+\lambda}+\frac{\left(\sum_{i\in I_{R}^{j}}g_{i}\right)^{2}}{\sum_{i\in I_{R}^{j}}h_{i}+\lambda}-\frac{\left(\sum_{i\in I^{j}}g_{i}\right)^{2}}{\sum_{i\in I^{j}}h_{i}+\lambda}\right],$ (28)
which is just the original split criterion applied to the data from one particular era. The era splitting criterion in its basic form is written as a mean average of the era-wise splitting criterion for each era of data.
@@
This criterion defined in equation (29) will have the highest score for the split which provides the greatest average decrease in impurity over all of the eras of data. Let us notice that this new criterion recovers equation 26 in the case where there is only one era of data. Therefore the era splitting criterion can be viewed as a generalization of the original splitting criterion. This criterion will be less-likely to favor splits which work really well in some eras but not in others. Splits are not considered which are undefined in any era. Potential split points must have data to the left and right of them to be considered as optimal. The reason for this choice is illustrated in figure 3. The plot on the left shows a data configuration where the original splitting criterion will choose a split which doesn’t induce an impurity decrease in any era of data by itself. On the other hand, the criterion in equation (29) chooses a split reducing impurity in both eras simultaneously.
A further generalization is made by way of a smooth max function, called the Boltzmann operator Asadi and Littman [2017]. The Boltzmann operator operates on an array of real numbers, and has one parameter, called $\alpha$. When $\alpha=0$, the Boltzmann operator recovers the arithmetic mean, when $\alpha=-\infty$ it recovers the $\min$, and when $\alpha=\infty$ it recovers the $\max$. It is known as a smooth max (min) function.
Definition 5 (The Boltzmann Operator).
Let $x_{i}$ for $i\in\{1,...,n\}$ be $n$ distinct real numbers, and $alpha\in[-\infty,\infty]$, then define the Boltzmann Operator $\mathcal{B}_{\alpha}$ as
$\mathcal{B}_{\alpha}(x_{1},...,x_{n})=\frac{\sum_{i=1}^{n}x_{i}\exp{(\alpha x_{i})}}{\sum_{i=1}^{n}\exp{(\alpha x_{i})}}.$ (30)
Proposition 2 (Limit of the Boltzmann Operator).
$\lim_{\alpha\to\infty}\mathcal{B}_{\alpha}(x)=\max(x)$ (31)
Proof.
Let
$\hat{x}=\max(x)$
and factor the $\hat{x}$ term from both the numerator and denominator of $\mathcal{B}_{\alpha}(x)$
$\mathcal{B}_{\alpha}(x)=\frac{\hat{x}e^{\alpha\hat{x}}+\sum_{x_{i}\neq\hat{x}}x_{i}e^{\alpha x_{i}}}{e^{\alpha\hat{x}}+\sum_{x_{i}\neq\hat{x}}e^{\alpha x_{i}}}.$ (32)
Now, divide numerator and denominator by $e^{\alpha\hat{x}}$,
$\mathcal{B}_{\alpha}(x)=\frac{\hat{x}+\sum_{x_{i}\neq\hat{x}}x_{i}e^{-\alpha(\hat{x}-x_{i})}}{1+\sum_{x_{i}\neq\hat{x}}e^{-\alpha(\hat{x}-x_{i})}}.$ (33)
The term $\hat{x}-x_{i}>0$ for all $x_{i}:x_{i}\neq\hat{x}$, which means that
$\lim_{\alpha\to\infty}e^{-\alpha(\hat{x}-x_{i})}=0,\forall x_{i}:x_{i}\neq\hat{x}.$ (34)
Thus, the limit of $\mathcal{B}_{\alpha}(x)$ is found by taking the limit of numerator and denominator, noticing that the sums of exponential functions all go to zero, yielding
$\lim_{\alpha\to\infty}\mathcal{B}_{\alpha}(x)=\frac{\hat{x}}{1}=\hat{x}.$ (35)
Proposition 2 for the limit and its proof can be given similarly to prove that $\lim_{\alpha\to-\infty}\mathcal{B}_{\alpha}(x)=\min(x)$. The final version of the era splitting criterion is defined next, which is understood as the smooth maximum (minimum) of the era-wise impurity reduction over all of the eras, and is a generalization of the original splitting criterion in Eq 26.
Definition 6 (Era Split Criterion).
The era split criterion $\mathcal{L}_{\text{era split}}^{\alpha}$ is defined by
$\mathcal{L}_{\text{era split}}^{\alpha}=\mathcal{B}_{\alpha}\left(\mathcal{L}_{\text{split}}^{1},...,\mathcal{L}_{\text{split}}^{M}\right).$ (36)
@@ 3 Directional Era Splitting
There is a subtle ambiguity that arises with equation 36, which requires the introduction of the concept of the direction implied by a split. A split implies a direction via the predicted values of the child nodes. Remember from proposition 1 that the predicted value at any node is the mean of all the dependent variables $y$ contained in the node. See figure 4. If the value of the left child node is larger than the value of the right child node, then the direction of this split is too the left. If the value of the right child node is larger, then the direction is to the right. Let the value of the left child node be $v_{l}$, and the value of the right child node be $v_{r}$, then without loss of generality, the direction of the split, $d_{\text{split}}$, is defined as the sign of the difference between the left child node value and the right child node value
$d_{\text{split}}=\text{sign}(v_{l}-v_{r}).$ (37)
In the original setting where all the training data is pooled together, there can be only one direction implied over the entire data set. The traditional splitting criterion (eq. 26) does not consider the direction of the split, since it is superfluous. It doesn’t matter whether the direction implied by the split is to the left or to the right, since there is only one era, just as long as the impurity is reduced. In era splitting, the split criterion is computed for each era separately, and then summarized with the Boltzmann operator. A problem that arises here is that there could be splits that imply conflicting directions from one era to the next. This is akin to single variable linear regression models having coefficients with opposite signs. We can describe the direction of a split in era $j$ by adapting equation 37 with era-wise descriptors, as follows.
@@
The directions implied by the split for each era of data results in an array of length $M$, $\{d_{\text{era split}}^{1},d_{\text{era split}}^{2},...,d_{\text{era split}}^{M}\}$.
Definition 7.
Directional era splitting, $\mathcal{D}_{\text{era split}}$, is the average agreement in the direction of a split over all $M$ eras.
$\mathcal{D}_{\text{era split}}=\frac{1}{M}\left|\sum_{j=1}^{M}d_{\text{era split}}^{j}\right|$ (39)
Each $d_{\text{era split}}^{j}$ will take a value of either $1$ or $-1$, depending on the direction of the split for that era. The sum is then bounded on the interval $[-M,M]$. By taking the absolute value and dividing my $M$, $\mathcal{D}_{\text{era split}}$ is then bounded on the interval $[0,1]$. At the low end, with a value of $\mathcal{D}_{\text{era split}}=0$, half of the directions go in one way while half go in the other way, the highest level of disagreement. At the high end, with a value of $1$, all the directions go the same way, the most agreement. When choosing the best split during tree growth, higher values of $\mathcal{D}_{\text{era split}}$ are preferred.
@@ 4 Theoretical Breakdown
This section expounds three propositions which are illustrative of the motivations behind era splitting. Here the focus is on era splitting (eq. 36), and not directional era splitting (eq. 39). The former being more anchored in traditional theory, and connected with the original criterion. The later is new and experimental, with impurity tied to the direction of the predictions in each era. Nevertheless, the following proposition 5 applies to both. This section contains all new, original content.
One problem with traditional splitting methods in the OOD setting is that the ”optimal” split can be degenerate in any or all of the eras (environments). The term, degenerate, indicates that the split doesn’t result in improved impurity in the child nodes compared to the parent for any particular era of the data set.
Definition 8 (Degenerate Splits).
A split is considered degenerate if $\mathcal{L}_{\text{split}}^{j}<=0$, or is undefined, for any era $j\in M$ for $M$ eras of training data. Where $\mathcal{L}_{\text{split}}^{j}$ is defined in equation (28). The expression is undefined if the split results in one completely empty child node.
This is a split that never would have been if the tree was grown on that era of data independently. The left side of figure 3 shows a simple example of a degenerate split with a toy data set of 4 data points. The original split criterion favors the split (red dashed line) which effectively separates era 1 data from era 2 data. In each era, this split does not induce any reduction in impurity. The era splitting criterion, on the other hand, favors a split (green dotted line) which reduces impurity in both eras simultaneously. Traditional splitting methods can induce splits that don’t split the data in some particular environment at all.
Proposition 3.
@@
Table 1: Data for proposition 3.
@@
This is one of the main conceptual motivations for the current research: it is possible that the original method misses, ignores, or misinterprets key environmental information. Era splitting addresses this problem in a very direct way: evaluate each split based on each environment separately. There is one configuration of era splitting which ensures the split points will always reduce impurity in every environment (era) of data simultaneously, never producing degenerate splits. This result is presented in proposition 4.
Proposition 4.
The era splitting criterion, as defined in equation (36), with Boltzmann alpha parameter approaching $\infty$- ∞ will never induce a degenerate split.